- Press Release
- Dec 8, 2022
NASA Space Station On-Orbit Status 30 Mar 2004
All ISS systems continue to function nominally, except those noted previously or below.
A major task to tackle for FE Alexander Kaleri today was the replacement of the failed Russian laptop 1 (LT1) with the borrowed U.S. laptop #6075 (an IBM ThinkPad 760XD), and to restore LT3 to its original function as payload computer, after its temporary service with the SM KTsP-1 (Service Module Central Post computer 1) as LT1’s replacement. More work remains to be done. [The swap-out required a number of scripted steps such as equipping the 760XD shell (SSC2) with the appropriate HDD (hard disk drive, #6059) and LT1 EtherLink card, switching SM control over to the secondary KTsP-2 and its controlling LT2, and installing #6075 at the Central Post (TsP). The procedures, supported by tagup with ground specialists, were not completed. To test the new LT1 configuration, the ground will need to switch back to the new primary LT1/KTsP-1, leaving it in operation if the test is successful. Currently, KTsP-2 remains “in charge” at the SM TsP.]
Correlated computer activity was also on CDR/SO Michael Foale’s schedule, to prepare and reconfigure a ThinkPad A31p NGL (Next Generation Laptop) from stowage as an additional workstation, by loading it with the new version 5.00 SSC software from the NGSD (Next Generation Support Disk), i.e., the transition load for the next Increment (Expedition 9). [As part of the activity, if time permitted, Foale was to perform troubleshooting on the shell of 760XD #6066, the stowed original SSC2 (station support computer 2) at the SM TsP which had failed to boot up properly with known good hard drive, and to replace it, if necessary, with the SSC7 (#6006) shell to become the new SSC2 Client at the TsP.]
The Flight Engineer continued his current round of monthly preventive maintenance on Russian segment air ventilation systems, today in the “Pirs” DC-1 docking compartment, where he cleaned the two PF1 & PF2 dust filters of its air duct system as well as the protective mesh screens of the V1 & V2 ventilator fans (last time done: 1/20).
Mike Foale tagged up with the BCAT (Binary Colloidal Alloy Test) payload developer to discuss yesterday’s experiment run. Additional photo imagery was obtained today with the Kodak 760 digital still camera for downlink. [Homogenization of the BCAT-3 samples yesterday went “well (and quickly!)”, according to first ground evaluation.]
A planned noise/vibration check by Alex Kaleri on the RPD (pressure differential regulator) #42, removed from the Elektron O2 generator on 2/13, was cancelled for today.
Mike Foale set up the joint US/Russian sound level meter (SLM) in the cabin for subsequent acoustic survey ops for ~ 1.5 hrs. Its recorded data were later transferred to the medical equipment computer (MEC). [These acoustic measurements are obtained once per month at 41 locations in the Lab, Node, Airlock, FGB, SM and DC-1 modules. The SLM gives instantaneous noise levels and their frequency spectra, which are transferred to the MEC laptop via an RS232 cable and later downlinked with regular CHeCS (crew health care systems) data dump or via OCA.]
In preparation for the upcoming R&R (removal & replacement) of the VD1 and VD2 fans of the Soyuz TMA-3/7S’s thermal control system, Kaleri first unstowed and relabeled the ZUP-86-II control/safety device for the fan system, which he had found on 3/9 in the FGB after some search, then prepared two spare fans from the FGB for the replacement. [During yesterday’s troubleshooting of the KhSA cooling/drying unit’s primary fan (VD1) in the Soyuz Descent Module (SA), which had failed on 10/18 during the 7S free flight, the FE bypassed the automatic switch for the backup fan with a jumper, but the fan still did not start. It was then found to be binding, i.e., unable to rotate freely.]
Sasha Kaleri performed the periodic IMS (inventory management system) medical equipment audit in the SM, comparing specific items against an uplinked list prepared with IMS data and sorting them out for disposal or stowage.
TVIS exercise data files were downloaded by the crew to the MEC, as is standard procedure whenever more than four sessions have been performed on the treadmill since last download.
Sasha also continued yesterday’s session of the Russian Uragan earth-imaging program, today focusing the Kodak DCS760 digital camera with 800-mm lens on new targets of nature and industry environment conditions. [They included the Persian Gulf coast, United Arab Emirates, Vanch River valley, glacier at the head of this valley all the way to the wide Alaysk valley, the Southern coast of Lake Issyk-Kul, the Kerch Strait, and the Kuban River flood plain.]
Sasha prepared the daily IMS inventory “delta” file for automated updating the IMS databases, and Mike held his periodic tagup with MCC-Houston IMS specialists via S/G1 (Space-to-Ground 1 audio).
Both crewmembers completed their daily physical exercise program of 2.5 hrs. each.
At 11:15am EST, the crew participated in a 20-min. educational PAO event with the Aerospace Academy for Engineering and Teacher Education at the Space Center Houston. Questions from the students from Alvin, Clear Creek, Friendswood, Galena Park, Galveston and Pasadena Independent School Districts in Texas, gathered at the event, had been uplinked beforehand.
During last night’s start of modes checkout of the S1 truss TRRJ (thermal radiator rotary joint), the Failure Detection, Isolation, and Recovery system (FDIR) was triggered by an undervoltage indication from the joint resolver (which keeps track of the RJ’s actual rotation angle). [Cause was found to be an inaccurate (i.e., too high) limit setting in the prepositioned load (PPL). The checkout is continuing and the limit values will be updated to be more realistic.]
Starting this morning at 3:00am EST and ending at about 9:00am, MCC-Houston and its support group in Moscow (HSG) performed another BCC (backup control center) activation test. As usual, it did not involve the ISS crew. [Purpose of the BCC checkout is to demonstrate procedural functionality and provide proficiency training for HSR (Houston Support Room) personnel. BCC utilizes the HSG command server and telemetry server. In test mode, command function is transferred to HSG (as BCC) for sending a single PPCP (preplanned command packet) with the primary command & control systems (CCS) “No Op Test” command to MCC-M/TsUP. TsUP confirms receipt of the PPCP, verifying telemetry by playing a data tape of previously recorded ACT (American Contingency Telemetry) for the HSR, upon which the latter cancels the PPCP. TsUP commanding of to the Russian segment (RS) via U.S. S-band as well as MCC-H-to-TsUP flow of status telemetry are not affected, but MCC-H has no real-time telemetry and command capability via Russian ground stations (RGS) during the periodic testing.]
Beginning at 7:30am today, during three orbits the periodic U.S. solar array efficiency test was performed, which assesses the degradation of the photovoltaics over time. Before the test, at 7:00am, the BGAs (beta gimbal assemblies) were moded to Autotrack (from fixed dual-angle) to improve power generation during the test, and afterwards returned to dual angle, with “night glider” drag reduction feathering. [For the test, the solar array output is shunted (switched) for ~10 min. such that the SSU (sequential shunt unit) circuit current can be recorded to determine solar array output capability. The amount of current shunted by the arrays will decrease slowly over time as they degrade in the orbital environment. The test is performed periodically (approximately twice a year) at a Beta angle of 0 plus/minus 5 deg. to allow engineers to track the performance over time and compare degradation trends against the a priori (expected) curve.]
Crew time will be scheduled for replacing the pulley Flexpacks that have failed in the RED (resistive exercise device) canisters #1011 & #1012. The task will take approximately three hours.
Today’s CEO (Crew Earth Observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were SE Asia smoke (Dynamic event. Regional smoke palls continue to make the news. Looking left into the Irrawady basin and right into the Mekong basin, both dry with significant biomass burning), Rangoon, Myanmar (looking slightly right of track on the far side (east side) of the Irrawady River delta. The city lies inland from the delta mouth), North China Plain air quality (winds from the Gobi have cleared the air in the Beijing region. Obliques are requested for comparison with images of days when air pollution is heavy), Tianjin, China (nadir pass over Beijing’s great port city), Yellow River Delta (this highly changeable delta is one of several major deltas for which the length of record of handheld imagery has revealed interesting details of landform evolution. Nadir pass), Bombay, India (looking a touch left), Iguazu, Argentina (one of the fastest growing cities in South America, Iguazu lies at the junction of Argentina, Brazil and Paraguay. Tourism related to the nearby Iguazu Falls [South America’s second tourist attraction after the ruins at Macchu Pichu], is now less important than trade for the local economy), Alps snowpack (Dynamic event. Panoramas of spring snowpack are requested), La Paz, Bolivia (nadir pass), Puerto Rico (city growth on the north coast is the main topic of interest, with impacts of urbanization on beaches and coral reefs as secondary interests), Pastures research, N Texas (detailed images of pasture color either side of the Red River [the major river in view and easily detected] were requested for studies of soil moisture/fertility. Shooting both sides of track near nadir), and St. Louis, Missouri (looking right of track at the junction of the Mississippi and Missouri Rivers).
CEO images can be viewed at the websites.
See also the website “Space Station Challenge” at
U.S. and Russian Segment Status (as of today, 2:12pm EST).
Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
- Elektron O2 generator is On. Vozdukh CO2 scrubber is On. U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
- SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.7; ppO2 (mmHg) — 159.5; ppCO2 (mmHg) — 3.7;
- SM Transfer Compartment: Pressure (mmHg) — 752; temperature (deg C) — 20.1.
- FGB Cabin: Pressure (mmHg) — 752; temperature (deg C) — 22.7.
- Node: Pressure (mmHg) — 750.89; temperature (deg C) — 22.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- U.S. Lab: Pressure (mmHg) — 753.0; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- Joint Airlock (Equip. Lock): Pressure (mmHg) — 753.2; temperature (deg C) — 23.1; shell heater temp (deg C) — 22.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
- (n/a = data not available)
- PMA-1: Shell heater temp (deg C) — 23.0
- PMA-2: Shell heater temp (deg C) — 20.3
Electrical Power Systems (EPS):
- Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in directed position (Blind mode, non solar-tracking, drag reduction-biased).
- SM batteries: Battery #8 is off-line; all other batteries (7) are in “Partial Charge” mode.
- FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); battery #3 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
- Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.
Command & Data Handling Systems:
- C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
- GNC-1 MDM is prime; GNC-2 is Backup.
- INT-2 is operating; INT-1 is Off.
- EXT-1 is On (primary), EXT-2 is Off.
- LA-1, LA-2 and LA-3 MDMs are all operating.
- PL-1 MDM is Off; PL-2 MDM is Operational.
- APS-1 (automated payload switch #1) and APS-2 are both On.
- SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
- SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22/03).
- FGB MDM-1 is Off (failed); MDM-2 is Operational.
- Total propellant load available: 3926 kg (8655 lb) as of 3/26/04; [SM(755) + FGB(2512) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).
Attitude Control Systems:
- 3 CMGs on-line (CMG-1 failed).
- State vector source — U.S. SIGI-1 (GPS)
- Attitude source — U.S. SIGI-1 (GPS)
- Angular rate source — RGA-1
- XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.
Communications & Tracking Systems:
- FGB MDM-1 is powered Off; FGB MDM-2 is operational.
- All other Russian communications & tracking systems are nominal.
- S-band is operating nominally (on string 2).
- Ku-band is operating nominally.
- Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
- Video subsystem operating nominally.
- HCOR (high-rate communications outage recorder) is operating nominally.
- SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
- MBS: KA (keep alive) power on both strings.
- MT: latched and mated at WS4.
- POA: KA power on both strings.
- RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.
ISS Orbit (as of this morning, 7:29am EST [= epoch]):
- Mean altitude — 365.8 km
- Apogee — 373.0km
- Perigee — 358.6 km
- Period — 91.86 min.
- Inclination (to Equator) — 51.6301 deg
- Eccentricity — 0.00107
- Solar Beta Angle — 18.2 (magnitude decreasing)
- Orbits per 24-hr. day — 15.67
- Mean altitude loss last 24 hours — 200 m
- Revolutions since FGB/Zarya launch (Nov. ’98) — 30596
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see